OpenStudy (anonymous):

Evaluate the expression for x = 4 and y = 5. 6(x² - y) - y²

OpenStudy (anonymous):

6(4^2-5)-5^2 6*4^2= 96 6*-5= -30 5^2 = -25 End up with 96-30-25 = 41

OpenStudy (anonymous):

Thank you, I actually figured it out while you were replying!

OpenStudy (anonymous):

Awesome :)